[Brahms-l] Shift report 20050619 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 19 2005 - 15:56:39 EDT
Shifter leader was JHL
Shifters were 


1. Beam:
   - Physics Beam for the entire shift: 8 hours
   - ZDC: 800 - 600/sec with 111 bunches.
   - Polarization B/Y = 47/50%

2. Data: 
   - FS at 2.3 deg. 1/1B
   - Runs 15656 - 15670
   - 2.2 M Trigger2 Events
   - No MRS data taken due to TPM2 Drift Voltage Problem

3. Problems:
   - HV 1458A/slot7 for TPM2 Drift Voltage gone bad
     (Need an access. Will be fixed tomorrow)
   - Old problems: 2 bad TDCs in H1, T3-3   

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 19 15:56:54 2005

This archive was generated by hypermail 2.1.8 : Sun Jun 19 2005 - 15:57:01 EDT